OpenStudy (anonymous):

HOW MUCH PURE ACID SHOULD BE MIXED WITH 5 GALLONS OF A 30% ACID SOLUTION IN ORDER TO GET A 50% ACID SOLUTION?

OpenStudy (amistre64):

p(1) + 5(.3) = (p+5)(.5)

OpenStudy (amistre64):

p + 1.5 = .5p + 2.5 .5p = 1 p = 1/.5 = 10/5 = 2 2 gallons

OpenStudy (anonymous):

thank you so much!